OpenStudy (anonymous):

2cos(2x)*sin(x+PI/4)-sin(2x)*cos(x+PI/4)=0 Solve.

OpenStudy (y2o2):

\[2\cos(2x)\sin(x+{\pi \over 4}) - \sin(2x)\cos(x+{\pi \over 4}) = 0\] ?

OpenStudy (y2o2):

did i type it right ?

OpenStudy (anonymous):

yes you did.